How many pounds of a certain alloy containing 35 percent must be melted with another alloy containing 65 percent silver to obtain 50 pounds of an alloy containing 50 percent silver?
Answer by mananth(16949)

percent ----------------quantity
35.00% ----------------x
65.00% ----------------50-x
Total 50.00%----------------50
35x+65(50-x)=50*50
35x+3250-65x=2500
35x-65x =2500-3250
-30x-=-750
/-30
x=25 pounds 35% alloy
25 pounds 65% alloy


25